OpenStudy (anonymous):

evaluate the integral ∫x^2 cos2x dx

sam (.sam.):

integration by parts

sam (.sam.):

\[\int\limits x^2 \cos (2 x) \, dx=\frac{1}{2} x^2 \sin (2 x)-\int\limits x \sin (2 x) \, dx\]

sam (.sam.):

integration by parts again \[\frac{1}{2} x^2 \sin (2 x)+\frac{1}{2} x \cos (2 x)-\frac{1}{2}\int\limits \cos (2 x) \, dx\]

sam (.sam.):

for cos2x , substitution u=2x , du=2dx \[\frac{1}{2} x^2 \sin (2 x)+\frac{1}{2} x \cos (2 x)-\frac{1}{4}\int\limits \cos (u) \, du\]

OpenStudy (anonymous):

actually the question jst says evaluate the integral nt intergration by parts

sam (.sam.):

integrate\[ x^2 \sin (x) \cos (x)-\frac{1}{4} \sin (2 x)+\frac{1}{2} x \cos (2 x)+c\]

sam (.sam.):

evaluate=integrate?

OpenStudy (anonymous):

thats jst what it said...-_-, cud u show the steps, please?

sam (.sam.):

those are the steps

OpenStudy (anonymous):

thanx :)

sam (.sam.):

welcome